Hi can you help me solve this equation?
3(2x-1)=5(3x-15)
hope you can help

Answer by venugopalramana(3286)

3(2x-1)=5(3x-15)
3*2x-3*1=5*3x-5*15
6x-3=15x-75
15x-75=6x-3
15x-6x=-3+75
9x=72
x=72/9=8